OpenStudy (anonymous):

Will medal if right Francine and Cheryl received equal scores on a test made up of multiple choice questions and an essay. Francine got 34 multiple choice questions correct and received 14 points for her essay. Cheryl got 30 multiple choice questions correct and received 22 points for her essay. How many points was each multiple choice question worth?

OpenStudy (anonymous):

34x+14 = 30x+22 Solve for X

OpenStudy (anonymous):

34x+14 = 30x+22

OpenStudy (anonymous):

34x+14 = 30x+22 yeah you have to solve for x

OpenStudy (anonymous):

I aggee

OpenStudy (anonymous):

34m + 14 = 30m + 22 4m = 8 m = 2

OpenStudy (anonymous):

K I did it I got 2 also after u gave me the equation I hate doing these word things